# Exact Solutions and Quantum Defect Theory for van der Waals Potentials in Ultracold Molecular Systems

### v3

@@ Abstract
In this paper, we have provided exact two-body solutions to the 2D and 3D Schrödinger equations with isotropic van der Waals potentials of the form ±$\pm 1/r^{6}$. Based on these solutions, we developed an analytical quantum defect theory (QDT) applicable to both quasi-2D and 3D geometries, and applied it to study the scattering properties and bound-state spectra of ultracold polar molecules confined in these geometries. Interestingly, we find that in the attractive (repulsive) van der Waals potential case, the short-range interaction can be effectively modeled by an infinite square barrier (finite square well), which leads to narrow and dense (broad and sparse) resonance structures in the quantum defect parameter. In the quasi-2D attractive case, shape resonances can appear in an ordered fashion across different partial waves, characterized by sharp phase jumps as the scattering energy is varied. Furthermore, the low-energy analytical expansions derived from QDT show excellent agreement with the exact numerical results, validating the accuracy and usefulness of our analytical approach in describing two-body physics governed by long-range van der Waals interactions.
@@ 2 Solutions of the Schrödinger equation
We analyze the radial Schrödinger equation with van der Waals potentials of the form ±$\pm C_{6}/r^{6}$ in two and three dimensions, governing the radial wave function $\overline{u}_{\overline{\epsilon}\overline{l}}(r)$, where $\overline{\epsilon}$ and $\overline{l}$ are the reduced energy and effective angular momentum, respectively:
$\left[\frac{d^{2}}{dr^{2}}-\frac{\overline{l}(\overline{l}+1)}{r^{2}}-\delta\frac{\beta_{6}^{4}}{r^{6}}+\overline{\epsilon}\right]\overline{u}_{\overline{\epsilon}\overline{l}}(r)=0.$ (1)
Here, $\beta_{6}=(2\mu C_{6}/\hbar^{2})^{1/4}$ defines the van der Waals length, with $\mu$ and $C_{6}$ being the reduced mass and interaction strength, respectively. The reduced energy is defined as $\overline{\epsilon}=2\mu\epsilon/\hbar^{2}=k^{2}$, where $\epsilon$ is the scattering energy and $k$ is the relative wave number. In 2D, $\overline{l}=m-1/2$, where $m$ is the azimuthal quantum number; in 3D, $\overline{l}=l$, where $l$ is the orbital angular momentum quantum number. The index $\delta$ distinguishes between repulsive ($\delta=+1$) and attractive ($\delta=-1$) potentials.
To derive analytical solutions, we first transform Eq. (1) into a dimensionless form using the following substitutions:
where $\Delta=\overline{\epsilon}L^{2}/8$ is the scaled energy and $\nu_{0}=(2\overline{l}+1)/4$. For the attractive van der Waals potential in three dimensions, the solution to Eq. (3) was obtained in Ref. [20] using the Neumann expansion method [30, 31, 32]. This method has also been extended to dipole interactions (±$\pm C_{3}/r^{3}$) in both 2D [21] and 3D [22]. This approach allows us to derive solutions in a unified manner for both attractive and repulsive van der Waals potentials in 2D and 3D, with the corresponding definitions of the scaling length $L$ and the effective angular momentum quantum number summarized in Table 1. We now present the explicit forms of the exact solutions to Eq. (1) in the following part of this section.
Attractive: $-C_{6}/r^{6}$ Repulsive: $+C_{6}/r^{6}$
@@ 2.1 Summary of the generalized Neumann expansion solutions
We find that there exists a pair of linearly independent and real solutions with energy-independent asymptotic behavior near the origin ($r\ll\beta_{6}$). These solutions can be expressed explicitly as
[…]
First, the functions $\overline{u}_{\overline{\epsilon}\overline{l}}^{1}(r)$ and $\overline{u}_{\overline{\epsilon}\overline{l}}^{2}(r)$ in Eqs. (47) form another pair of linearly independent solutions, expressed as generalized Neumann expansions:
@@
with $j$ being a positive integer and the function $c_{j}$ defined as
@@
Second, the normalization coefficients in Eqs. (47) are given by
@@
Finally, the index parameter $\nu$ appearing in Eqs. (821) is a root of the characteristic function
where
$\overline{Q}(\nu)=\frac{Q(\nu)}{(\nu+1)[(\nu+1)^{2}-\nu_{0}^{2}]}.$ (24)
The solution $\nu$ to the equation $\Lambda_{\overline{l}}(\nu,\Delta^{2})=0$ can be either real or complex, depending on the scattering energy and angular momentum. The determination of $\nu$ is a key step in constructing the exact solution. A detailed analysis of the energy dependence of $\nu$, as well as its low-energy expansion for different partial waves, will be presented in Sec. 2.3.
@@ 2.2 Asymptotic behavior
where the partial wave index $\widetilde{l}$ equals $l$ in 3D and $m$ in 2D. For negative energy $\overline{\epsilon}=k^{2}=-\kappa^{2}<0$, the large-$r$ asymptotic behavior becomes
@@
These relations are independent of the energy $\overline{\epsilon}$ and angular momentum $\overline{l}$, and hold universally for all cases listed in Table 1. We have verified them in our calculations as a nontrivial consistency check of the solution structure. The explicit results for the $W^{\pm}$ and $Z^{\pm}$ matrices are provided in A.
2.3 Structure of $\nu$
In Fig. 1, we show the energy dependence of $\nu$ for the $s$-, $p$-, $d$-, and $f$-waves in 2D (a1–d1) and 3D (a2–d2). In 3D (a2–d2), each partial wave has a well-defined critical energy beyond which $\nu$ becomes complex, i.e., when $|\Delta^{2}|>|\Delta_{c}^{2}|$. In contrast, for 2D systems, $\nu$ can become purely imaginary even for the lowest partial waves. For general values of $l$ or $m$, the real part of $\nu$ tends to plateau as $|\Delta^{2}|$ increases beyond the corresponding critical energy.
2D 3D
Partial Wave Repulsive Attractive Repulsive Attractive
$s$ $-0.0169$ 00 $-0.00933$ $0.00932$
$p$ 00 $0.0253$ $-0.0218$ $0.0217$
$d$ $-0.103$ 00 $-0.209$ $0.185$
$f$ $-1.60$ $1.59$ $-1.17$ $2.50$
which reduces to $\nu_{0}$ when $\Delta\rightarrow 0$. Note that the expansion diverges for $\nu_{0}=0,1/2,1,2$, corresponding to $m=0,1,2,4$ in 2D systems. For these special cases, the low-energy behavior can instead be expressed as:
@@ 3 Quantum defect theory of the van der Waals potentials in quasi-2D and 3D
For ultracold and dilute molecular gases, the average distance between two molecules is much larger than the van der Waals length and then we usually can model the two-body interaction by the contact interaction associated with the scattering length which contains the information of short range potentials [34]. Either when one wants to have more accurate two-body scattering properties or when the molecular gases are not enough cold and dilute, we need to consider the van der Waals type shorter range potential, which can be induced from two polar molecules tuning by electric fields [18, 19]. In these experiments [18, 19], two 40K87Rb molecules prepared in the $|N,m_{N}\rangle=|1,0\rangle$ state, where $N$ is the rotational angular momentum and $m_{N}$ is its projection along the axis of electric field ${\bf E}$. Starting from this two-molecular state $|1,0\rangle\otimes|1,0\rangle$ at weak electric field ${\bf E}$, they increased the strength of electric field $|{\bf E}|$ to across two crossing energy points, where the two molecular states are $|0,0\rangle\otimes|2,\pm 1\rangle$ and $|0,0\rangle\otimes|2,0\rangle$, respectively. Near the crossing points, the two molecular states are coupled by the dipole interaction which weakly opens the energy gap via a second order perturbation process. Thus the effective attractive (repulsive) van der Waals barrier in the energy curve of these two molecules associated with state $|0,0\rangle\otimes|2,\pm 1\rangle$ ($|0,0\rangle\otimes|2,0\rangle$) emerges. This new technique for inducing the effective van der Waals potentials both be realized in quasi-2D [18] and 3D [19]. Next, we will construct the quantum defect theories for these systems.
@@ 3.1 Quantum defect parameter $K_{c}$
[…]
@@
[…]
@@
Therefore, whether one aims to determine phase shifts via Eqs. (40)–(41) or extract binding energies from Eq. (45), the quantum defect parameter $K_{c}^{(\pm)}$ must be specified. Moreover, due to the energy- and angular-momentum-insensitive nature of the wave functions in the short-range region $r\ll\beta_{6}$ (as shown in Eqs. (25)–(28)), the quantum defect parameter $K_{c}^{(\pm)}$ is also approximately independent of energy and partial wave at short distances. To illustrate this property, we divide the total potential into a short-range part and a long-range tail using a boundary located at $r=r_{0}$. The long-range tail is governed by the van der Waals potential ±$\pm C_{6}/r^{6}$, while the short-range component is modeled using idealized potentials: a finite square well for the repulsive case, and an infinite square barrier for the attractive case. These model potentials are shown schematically in Eqs. (46)–(47) and illustrated in Fig. 2(a,c), respectively.
@@
In Fig. 2, we show the behavior of the quantum defect parameter $K_{c}^{(\pm)}$ as a function of short-range potential parameters. The black solid lines correspond to 3D geometries, while the blue dashed lines represent 2D results. For the attractive cases [Fig. 2(b)], the density of shape resonances increases rapidly as the short-range cutoff $r_{0}$ approaches zero. In contrast, in the repulsive cases [Figs. 2(d–f)], the resonances are significantly sparser and much narrower in width. In fact, $K_{c}^{(+)}$ remains close to zero over most of the parameter space, and no prominent resonance emerges even as $r_{0}\rightarrow 0$, given a fixed potential depth $V_{0}$. This behavior can be attributed to the presence of a large centrifugal and repulsive van der Waals barrier, which effectively suppresses the influence of the short-range region on the scattering wave function. Consequently, for the repulsive case, we will focus on the pure repulsive limit ($r_{0}\rightarrow 0$), where $K_{c}^{(+)}=0$. For the attractive case, however, both scattering and bound-state properties will be analyzed across shape resonances, where $K_{c}^{(-)}$ can be tuned continuously from +$\infty$+ ∞ to $\infty$- ∞.
@@ 3.2 Scattering properties and bound states
[…]
In Fig. 3, we plot the phase shifts and cross sections for the first four partial waves in 2D and 3D. The phase shifts start from zero and increase with energy. Notably, each time the phase shift crosses an integer multiple of $\pi$, the corresponding cross section exhibits a dip due to destructive interference (see, for instance, the black curve in panel (b)).
At low energies ($k_{s}\ll 1$), the scattering is dominated by the $s$-wave contribution, while higher partial waves become relevant at larger energies. In the limit $k\rightarrow 0$, the cross section converges to
$\sigma_{l=0}^{3D}\rightarrow\pi\left[\frac{\Gamma(3/4)}{\Gamma(5/4)}\right]^{2},$
whereas in 2D it diverges logarithmically. This divergence is a well-known feature of 2D scattering, arising due to the long-range logarithmic behavior of the 2D Green’s function and the lack of a centrifugal barrier in the lowest channel.
@@
[…]
@@
where $B=2\ln k_{s}+3\gamma-4\ln 2$. From Fig. 5, we observe that the shallowest bound state emerges only in the $s$-wave channel as $K_{c}^{(-)}\rightarrow+\infty$, while deeper bound states appear sequentially across all partial-wave branches as $K_{c}^{(-)}$ decreases. All numerical results (dots) show excellent agreement with the low-energy analytical predictions (lines) in the weak-binding regime.

### v4

@@ Abstract
In this paper, we have provided exact two-body solutions to the 2D and 3D Schrödinger equations with isotropic van der Waals potentials of the form $\pm 1/r^{6}$. Based on these solutions, we developed an analytical quantum defect theory (QDT) applicable to both quasi-2D and 3D geometries, and applied it to study the scattering properties and bound-state spectra of ultracold polar molecules confined in these geometries. Interestingly, we find that in the attractive (repulsive) van der Waals potential case, the short-range interaction can be effectively modeled by an infinite square barrier (finite square well), which leads to narrow and dense (broad and sparse) resonance structures in the quantum defect parameter. In the quasi-2D attractive case, shape resonances can appear in an ordered fashion across different partial waves, characterized by sharp phase jumps as the scattering energy is varied. Furthermore, the low-energy analytical expansions derived from QDT show excellent agreement with the exact numerical results, validating the accuracy and usefulness of our analytical approach in describing two-body physics governed by long-range van der Waals interactions.
@@ 2 Solutions of the Schrödinger equation
We analyze the radial Schrödinger equation with van der Waals potentials of the form $\pm C_{6}/r^{6}$ in two and three dimensions, governing the radial wave function $\overline{u}_{\overline{\epsilon}\overline{l}}(r)$ (see the A for the detailed derivation):
$\left[\frac{d^{2}}{dr^{2}}-\frac{\overline{l}(\overline{l}+1)}{r^{2}}-\delta\frac{\beta_{6}^{4}}{r^{6}}+\overline{\epsilon}\right]\overline{u}_{\overline{\epsilon}\overline{l}}(r)=0,$ (1)
where $\overline{\epsilon}$ and $\overline{l}$ are the reduced energy and effective angular momentum, respectively. Here, $\beta_{6}=(2\mu C_{6}/\hbar^{2})^{1/4}$ defines the van der Waals length, with $\mu$ and $C_{6}$ being the reduced mass and interaction strength, respectively. The reduced energy is defined as $\overline{\epsilon}=2\mu\epsilon/\hbar^{2}=k^{2}$, where $\epsilon$ is the scattering energy and $k$ is the relative wave number. In 2D, $\overline{l}=m-1/2$, where $m$ is the azimuthal quantum number; in 3D, $\overline{l}=l$, where $l$ is the orbital angular momentum quantum number. The index $\delta$ distinguishes between repulsive ($\delta=+1$) and attractive ($\delta=-1$) potentials.
Attractive: $C_{6}/r^{6}$ Repulsive: $C_{6}/r^{6}$
@@
In this work, we follow the approach first developed by Gao in his exact solutions to the 3D attractive van der Waals potential and the repulsive dipole-dipole interaction [21, 22]. To derive analytical solutions, we begin by transforming Eq. (1) into a dimensionless form using the following substitutions:
where $\Delta=\overline{\epsilon}L^{2}/8$ is the scaled energy and $\nu_{0}=(2\overline{l}+1)/4$. For the attractive van der Waals potential in three dimensions, the solution to Eq. (3) was obtained in Ref. [21] using the Neumann expansion method [31, 32, 33]. This method has also been extended to dipole interactions ($\pm C_{3}/r^{3}$) in both 2D [23]. This approach allows us to derive solutions in a unified manner for both attractive and repulsive van der Waals potentials in 2D and 3D, with the corresponding definitions of the length scale $L$ and the effective angular momentum quantum number summarized in Table 1. It is worth noting that our Eq. (2) and Eq. (3) are formally identical to Eqs. (32-34) in Gao’s treatment of the 3D attractive van der Waals potential [21]. However, the key difference lies in the definition of the length scale $L$, as summarized in Table 1, and the associated scaled energy parameter $\Delta$. These distinctions lead to a different set of exact solutions and corresponding asymptotic behaviors. We now present the explicit forms of the exact solutions to Eq. (1) in the following part of this section. The results obtained in Secs. 2.1 and 2.2 are valid for all cases listed in Table 1.
@@ 2.1 Summary of the generalized Neumann expansion solutions
By solving the second-order differential equation Eq. (3) and substituting the results into Eq. (2), we obtain a pair of linearly independent solutions of Eq. (1), denoted as $\overline{u}_{\overline{\epsilon}\,\overline{\ell}}^{(1)}(r)$ and $\overline{u}_{\overline{\epsilon}\,\overline{\ell}}^{(2)}(r)$, which can be expressed in the form of generalized Neumann expansions:
@@
with $j$ being a positive integer, $\nu_{0}=(2\bar{l}+1)/4$ and the function $c_{j}$ defined as
@@
The index parameter $\nu$ appearing in Eqs. (49) is a root of the characteristic function
where
$\overline{Q}(\nu)=\frac{Q(\nu)}{(\nu+1)[(\nu+1)^{2}-\nu_{0}^{2}]}.$ (11)
The solution $\nu$ to the equation $\Lambda_{\overline{l}}(\nu,\Delta^{2})=0$ can be either real or complex, depending on the scattering energy and angular momentum. The determination of $\nu$ is a key step in constructing the exact solution. A detailed analysis of the energy dependence of $\nu$, as well as its low-energy expansion for different partial waves, will be presented in Sec. C.
We find that there exists a pair of linearly independent and real solutions with energy-independent asymptotic behavior near the origin ($r\ll\beta_{6}$). These solutions for the replusive potentials can be expressed explicitly as
$u_{\overline{\epsilon}\overline{l}}^{+1}(r)$ $\!=\!$ $\widetilde{X}^{-1}\left[i\overline{u}_{\overline{\epsilon}\overline{l}}^{1}(r)-\overline{u}_{\overline{\epsilon}\overline{l}}^{2}(r)\right],$ (12)
$u_{\overline{\epsilon}\overline{l}}^{+2}(r)$ $\!=\!$ $\widetilde{Y}^{-1}\overline{u}_{\overline{\epsilon}\overline{l}}^{1}(r),$ (13)
and for the attactive potentials, we have
$u_{\overline{\epsilon}\overline{l}}^{-1}(r)$ $\!=\!$ $\left(\alpha^{2}+\beta^{2}\right)^{-1}\left[\alpha\overline{u}_{\overline{\epsilon}\overline{l}}^{1}(r)-\beta\overline{u}_{\overline{\epsilon}\overline{l}}^{2}(r)\right],$ (14)
$u_{\overline{\epsilon}\overline{l}}^{-2}(r)$ $\!=\!$ $\left(\alpha^{2}+\beta^{2}\right)^{-1}\left[\beta\overline{u}_{\overline{\epsilon}\overline{l}}^{1}(r)+\alpha\overline{u}_{\overline{\epsilon}\overline{l}}^{2}(r)\right].$ (15)
The normalization coefficients in Eqs. (1215) are given by
@@ 2.2 Asymptotic behavior
where the index $\widetilde{l}$ can in principle be any integer. However, to maintain consistency with the standard asymptotic form of the scattering wavefunction at large distances, we set $\widetilde{l}=l$ in 3D and $\widetilde{l}=m$ in 2D. For negative energy $\overline{\epsilon}=k^{2}=-\kappa^{2}<0$, the large-$r$ asymptotic behavior becomes
@@
These relations are independent of the energy $\overline{\epsilon}$ and angular momentum $\overline{l}$, and hold universally for all cases listed in Table 1. Physically, this reflects the conservation of probability current in quantum mechanics, as the Wronskian is proportional to the flux carried by the wavefunctions. In scattering problems, this invariance ensures that boundary-matching conditions preserve the normalization and phase coherence of asymptotic solutions. We have verified them in our calculations as a nontrivial consistency check of the solution structure. The explicit results for the $W^{\pm}$ and $Z^{\pm}$ matrices are provided in B.
@@ 3 Quantum defect theory of the van der Waals potentials in quasi-2D and 3D
For ultracold and dilute molecular gases, the average distance between two molecules is much larger than the van der Waals length and then we usually can model the two-body interaction by the contact interaction associated with the scattering length which contains the information of short range potentials [35]. Either when one wants to have more accurate two-body scattering properties or when the molecular gases are not enough cold and dilute, we need to consider the van der Waals type shorter range potential, which can be induced from two polar molecules tuning by electric fields [18, 19]. In these experiments [18, 19], two 40K87Rb molecules prepared in the the rovibrational state $|\nu=0,N=1,m_{N}=0\rangle=|1,0\rangle$ of the singlet electronic potential $X^{1}\Sigma$ [20], where $\nu$ is the vibrational quantum number, $N$ is the rotational angular momentum and $m_{N}$ is its projection along the axis of electric field $\bf E$}. Starting from this two-molecular state $|1,0\rangle\otimes|1,0\rangle$ at weak electric field $\bf E$}, they increased the strength of electric field $|{\bf E}|$ to across two crossing energy points, where the two molecular states are $|0,0\rangle\otimes|2,\pm 1\rangle$ and $|0,0\rangle\otimes|2,0\rangle$, respectively. Near the crossing points, the two molecular states are coupled by the dipole interaction which weakly opens the energy gap via a second order perturbation process. Thus the effective attractive (repulsive) van der Waals barrier in the energy curve of these two molecules associated with state $|0,0\rangle\otimes|2,\pm 1\rangle$ ($|0,0\rangle\otimes|2,0\rangle$) emerges. This new technique for inducing the effective van der Waals potentials both be realized in quasi-2D [18] and 3D [19]. Next, we will construct the quantum defect theories for these systems.
@@ 3.1 Quantum defect parameter $K_{c}$
$\mbox{3D:}\quad\tan\delta_{l}^{(\pm)}$ $=$ $\frac{K_{c}^{(\pm)}Z_{22}^{(\pm)}-Z_{12}^{(\pm)}}{Z_{11}^{(\pm)}-K_{c}^{(\pm)}Z_{21}^{(\pm)}}$ (35)
$\mbox{2D:}\quad\tan\delta_{m}^{(\pm)}$ $=$ $\frac{K_{c}^{(\pm)}(Z_{21}^{(\pm)}+Z_{22}^{(\pm)})-(Z_{11}^{(\pm)}+Z_{12}^{(\pm)})}{(Z_{11}^{(\pm)}-Z_{12}^{(\pm)})-K_{c}^{(\pm)}(Z_{21}^{(\pm)}-Z_{22}^{(\pm)})}$ (36)
@@
$\mbox{3D:}\quad\sigma_{l}^{(\pm)}$ $=$ $4\pi(2l+1)\frac{\sin^{2}\delta_{l}^{(\pm)}}{k^{2}},$ (37)
$\mbox{2D:}\quad\sigma_{m}^{(\pm)}$ $=$ $4\frac{\sin^{2}\delta_{m}^{(\pm)}}{k}.$ (38)
@@
Therefore, whether one aims to determine phase shifts via Eqs. (35)–(36) or extract binding energies from Eq. (40), the quantum defect parameter $K_{c}^{(\pm)}$ must be specified. Moreover, in the short-range region $r\ll\beta_{6}$, the total energy $\overline{\epsilon}$ of the particle is typically much smaller than the local potential energy. As a result, the wavefunctions in this region become effectively insensitive to both energy and angular momentum, as demonstrated in Eqs. (25)–(28). Since the quantum defect parameter $K_{c}^{(\pm)}$ is determined by the logarithmic derivative of the short-range solutions, it also becomes approximately independent of energy and partial wave at short distances. To illustrate this property, we divide the total potential into a short-range part and a long-range tail using a boundary located at $r=r_{0}$. The long-range tail is governed by the van der Waals potential $\pm C_{6}/r^{6}$, while the short-range component is modeled using idealized potentials: a finite square well for the repulsive case, and an infinite square barrier for the attractive case. These model potentials are shown schematically in Eqs. (41)–(42) and illustrated in Fig. 1(a,c), respectively.
@@
In Fig. 1, we show the behavior of the quantum defect parameter $K_{c}^{(\pm)}$ as a function of short-range potential parameters. The black solid lines correspond to 3D geometries, while the blue dashed lines represent 2D results. For the attractive cases [Fig. 1(b)], the density of shape resonances increases rapidly as the short-range cutoff $r_{0}$ approaches zero. In contrast, in the repulsive cases [Figs. 1(d–f)], the resonances are significantly sparser and much narrower in width. In fact, $K_{c}^{(+)}$ remains close to zero over most of the parameter space, and no prominent resonance emerges even as $r_{0}\rightarrow 0$, given a fixed potential depth $V_{0}$. This behavior can be attributed to the presence of a large centrifugal and repulsive van der Waals barrier, which effectively suppresses the influence of the short-range region on the scattering wave function. Consequently, for the repulsive case, we will focus on the pure repulsive limit ($r_{0}\rightarrow 0$), where $K_{c}^{(+)}=0$. For the attractive case, however, both scattering and bound-state properties will be analyzed across shape resonances, where $K_{c}^{(-)}$ can be tuned continuously from $\infty$ to -$\infty$.
@@ 3.2 Scattering properties and bound states
$\mbox{3D:}\quad\tan\delta_{l}^{+}$ $=$ $-\frac{Z_{12}^{(+)}}{Z_{11}^{(+)}},$ (45)
$\mbox{2D:}\quad\tan\delta_{m}^{+}$ $=$ $-\frac{Z_{11}^{(+)}+Z_{12}^{(+)}}{Z_{11}^{(+)}-Z_{12}^{(+)}}.$ (46)
In Fig. 2, we plot the phase shifts and cross sections for the first four partial waves in 2D and 3D. The phase shifts start from zero and increase with the scaled momentum $k_{s}=k\beta_{6}$. Notably, each time the phase shift crosses an integer multiple of $\pi$, the corresponding cross section exhibits a dip due to destructive interference (see, for instance, the black curve in panel (b)).
At low energies ($k_{s}\ll 1$), the scattering is dominated by the $s$-wave contribution, while higher partial waves become relevant only at elevated energies. In the 3D limit $k\rightarrow 0$, the total scattering cross section asymptotically approaches a constant:
$\sigma_{\mathrm{tot}}^{\mathrm{3D}}=\sigma_{l=0}^{\mathrm{3D}}+\sigma_{l=1}^{\mathrm{3D}}+\sigma_{l=2}^{\mathrm{3D}}+\sigma_{l=3}^{\mathrm{3D}}+\sum_{j\geq 4}\sigma_{j}^{\mathrm{3D}}\approx\sigma_{l=0}^{\mathrm{3D}}\rightarrow\pi\left[\frac{\Gamma(3/4)}{\Gamma(5/4)}\right]^{2},$
where the low-energy expansions of the partial-wave contributions are given by
$\sigma_{l=0}^{\mathrm{3D}}$ $=\frac{\pi\,\Gamma\left(3/4\right)^{2}}{\Gamma\left(5/4\right)^{2}}-\frac{\pi\,\Gamma\left(3/4\right)^{4}}{12\,\Gamma\left(5/4\right)^{4}}\,k_{s}^{2},$ (47)
$\sigma_{l=1}^{\mathrm{3D}}$ $=\frac{\pi\,\Gamma\left(1/4\right)^{2}}{48\,\Gamma\left(7/4\right)^{2}}\,k_{s}^{4},$ (48)
$\sigma_{l=2}^{\mathrm{3D}}$ $=\frac{4\pi^{3}}{19845}\,k_{s}^{6},$ (49)
$\sigma_{l=3}^{\mathrm{3D}}$ $=\frac{4\pi^{3}}{1715175}\,k_{s}^{6}.$ (50)
In contrast, in the 2D case, the total cross section exhibits a logarithmic divergence as $k\rightarrow 0$. This divergence is a well-known feature of 2D scattering, arising due to the long-range logarithmic behavior of the 2D Green’s function and the lack of a centrifugal barrier in the lowest channel.
@@
For the attractive case in 2D, we investigate both scattering properties and bound-state spectra for the first four partial waves, as shown in Figs. 3 and 4. To illustrate the variation across resonance, we choose four representative values of the quantum defect parameter $K_{c}^{(-)}=10,\,0.1,\,-0.1,\,-10$ in Fig. 3.
@@
where $B=2\ln{\color[rgb]{0,0,0}|k_{s}|}+3\gamma-4\ln 2$. From Fig. 4, we observe that the shallowest bound state emerges only in the $s$-wave channel as $K_{c}^{(-)}\rightarrow{\color[rgb]{0,0,0}0^{-}}$, while deeper bound states appear sequentially across all partial-wave branches as $|K_{c}^{(-)}|$ increases. All numerical results (dots) show excellent agreement with the low-energy analytical predictions (lines) in the weak-binding regime.
For attractive interactions in 2D, the QDT parameter $K_{c}^{(-)}$ introduced in Eqs. (59)-(62) is directly related to the 2D scattering length $a_{2D}$. Specifically, by comparing the low-energy behavior of the $m=0$ phase shift,
$\tan\delta_{m=0}=\frac{\pi}{2\ln(ka_{2D})},\qquad k\rightarrow 0.$ (67)
Comparing this behavior to that in Eq. (59), one can immediately obtain the relations
$a_{2D}=e^{-\frac{\pi}{4K_{c}^{(-)}}}\bar{\beta}_{6}$ (68)
or
$K_{c}^{(-)}=-\frac{\pi}{4\ln(a_{2D}/\bar{\beta}_{6})}$ (69)
with $\bar{\beta}_{6}=e^{3\gamma/2}\beta_{6}/4$. Substituting this into Eq. (59), the $s$-wave phase shift can be expressed in terms of $a_{2D}$ as
$\tan\delta_{m=0}^{-}=\frac{\pi}{2\ln(ka_{2D})}-\frac{\pi^{2}+16\ln^{2}(a_{2D}/\bar{\beta}_{6})}{32\ln^{2}(ka_{2D})}\,k_{s}^{2}.$ (70)
Likewise, the low-energy bound state energy can also be expressed via $a_{2D}$. In the limit $a_{2D}\rightarrow\infty$, the $s$-wave binding energy reads
$E_{b}=-\frac{\hbar^{2}\kappa^{2}}{m},~~~\kappa a_{2D}\simeq 1-\left[\frac{1}{\pi}\ln^{2}\left(\frac{a_{2D}}{\bar{\beta}_{6}}\right)+\frac{\pi}{16}\right]\left(\frac{\beta_{6}}{a_{2D}}\right)^{2}.$ (71)